risolvi le seguenti equazioni.
-12x - (8x+4) + 10x = -2(x+1)
5 + 11x -2(2x+4) -3x = 10x + 9
3(x-2)+5(x+1)=-2(x+4)+5

Answer by checkley79(3341)

-12x - (8x+4) + 10x = -2(x+1)
-12x-8x-4+10x=-2x-2
-10x+2x=-2+4
-8x=2
x=2/-8=- 1/4 ans.
Proof:
-12(-1/4)-(8(-1/4)+4)+10(-1/4)=-2((-1/4)+1)
3-(-2+4)-10/4=-2(3/4)
3-2-2.5=-1.5
-1.5=-1.5
-------------------------
5 + 11x -2(2x+4) -3x = 10x + 9
5+11x-4x-8-3x=10x+9
4x-10x=9+3
-6x=12
x=12/-6
x=-2 ans.
Proof:
5+11*-2-2(2(-2)+4))-3(-2)=10(-2)+9
5-22-2(-4+4)+6=-20+9
5-22-2*0+6=-20+9
-11=-11
-----------------------------------------
3(x-2)+5(x+1)=-2(x+4)+5
3x-6+5x+5=-2x-8+5
8x+2x=-3+1
10x=-2
x=-2/10
x=-1/5 ans.
Proof:
3(-1/5-2)+5(-1/5+1)=-2(-1/5+4)+5
3(-11/5)+5(4/5)=-2(19/5)+5
-33/5+4=-38/5+5
-33/5+20/5=-38/5+25/5
-13/5=-13/5